# MILR: Improving Multimodal Image Generation via Test-Time Latent Reasoning

### v2

@@ 1 Introduction
Text-guided image generation is the task of synthesizing an image conditioned on a given text instruction. In recent years, the field has witnessed transformative progress: moving from generative adversarial models (goodfellow2014gan) to autoregressive and diffusion approaches (sun2024LlamaGen; chen2025janusprounifiedmultimodalunderstanding; esser2024sd3-med; flux2024). However, traditional models are limited in generating images in a single-shot fashion and thus are unable to resolve potential defects (huang2025t2icompbench++; niu2025wise). Inspired by the success of reasoning-augmented LLMs—such as OpenAI o1 (openai2024openaio1card) and DeepSeek-R1 (deepseekai2025deepseekr1incentivizingreasoningcapability)—that can reflect on and refine their thoughts, and answer accordingly, recent works have attempted to endow image generation models with reasoning ability (guo2025can; fang2025got; zhang2025reasongenr1cotautoregressiveimage; wu2025repromptreasoningaugmentedrepromptingtexttoimage; chern2025thinkinggeneratedimages).
Reasoning-augmented image generation models typically perform reasoning in two spaces: language and image spaces. Language reasoning primarily involves refining instructions to make them more comprehensible to the model (wu2025repromptreasoningaugmentedrepromptingtexttoimage; li2025reflectdit), and image reasoning iterates on the generation guided by a quality metric (guo2025can; zhuo2025reflectionflow). Early works implement reasoning either in the language or in the image space, lacking a mechanism for synergistic reasoning across the two spaces. To fill the gap, recent works resort to unified multimodal understanding and generation (MUG; jiang2025t2i-r1; duan2025got-r1; zhang2025reasongenr1cotautoregressiveimage), which natively supports language reasoning before generating images (referred to as multimodal image generation), facilitating cross-modal image-text reasoning. Despite the success, these approaches require carefully curated reasoning data and depend on model fine-tuning, rendering them complex and costly to develop in practice.
@@
Reasoning with MILR involves finding the best vector representations of image and text tokens that lead to improved image quality. We implement it using the policy gradient method (williams1992simple), where, at test time, the reward is computed by scoring the compatibility between the generated image and the given instruction. Crucially, gradients are only back-propagated to the cross-modal latent representations (i.e., the intermediate model outputs) obtained from the forward pass (see Figure 1), without altering any model parameters, and thus making MILR a test-time latent reasoning method.
We evaluate MILR on three widely-used image generation benchmarks: GenEval (ghosh2023geneval), T2I-CompBench (huang2023t2icompbench), and WISE (niu2025wise). MILR established new state-of-the-art results across all benchmarks. Notably, it achieves an overall score of 0.95 on GenEval, matching the best training-based model and outperforming the best test-time-scaling method by 4.4%. On the more challenging WISE benchmark, MILR obtains an overall score of 0.63, surpassing the strongest baseline model by 16.7%. Our further analysis reveals that the best performance of MILR is driven by its ability to perform joint image-text reasoning in a unified latent space. This unique advantage also makes MILR successful in instructions that involve challenging cultural and temporal reasoning.
@@ 2 Related Works
Remarkable progress has been achieved in reasoning-augmented LLMs (wei2022chain; deepseekai2025deepseekr1incentivizingreasoningcapability), but effectively integrating reasoning into conventional text-guided image generation models remains a challenge (sun2024LlamaGen; chen2025janusprounifiedmultimodalunderstanding; betker2023dalle; esser2024sd3-med). With the development of unified multimodal understanding and generation that can generate images and text in an interleaved manner, many works have explored reasoning for text-guided image generation. The predominant paradigm uses classical causal language modeling to unify image and text generation as next token prediction, enabling the model to make plans via chain-of-thought before generating images (fang2025got; deng2025bagel; xiao2025mindomni; cai2025IGTR; guo2025can; jiang2025t2i-r1; duan2025got-r1). Another line of research focuses on reasoning over images via the test-time scaling strategy (li2025reflectdit; zhuo2025reflectionflow). These works typically rely on an external critic model to provide feedback on further improvement. Unlike all these methods that explicitly perform reasoning over raw images and text, we use test-time optimization to refine the latent representations of image and text tokens, leading to a unified cross-modal latent reasoning method.
@@
Differently from explicit reasoning via a chain of thoughts (wei2022chain), latent reasoning refers to an implicit reasoning mechanism applied to the latent states (e.g., intermediate outputs) of the model. As in Transformer-based models, latent reasoning is typically implemented as spatial and temporal recurrences. Spatial recurrences implicitly deepen the model by iterating over the latent states between the Transformer layers (hao2024training; cheng2024compressed; zhang2025lightthinker; shen2025codi), while temporal recurrences refine the latent states through iterations across input tokens (dao2024transformers; geiping2025scaling). Under the recurrences is the idea of scaling up test-time computation for inference; however, this requires the recurrent modules to be pre-trained. In a similar vein to those, we iterate over the unified image-text latent states only at test time, without introducing and updating any model parameters.
@@
Reinforcement learning (RL) has been the key to eliciting the reasoning ability of large language models (deepseekai2025deepseekr1incentivizingreasoningcapability; openai2024openaio1card). Inspired by the success of GRPO, which is an improved RL algorithm over PPO (schulman2017ppo) and is used in DeepSeek-R1 (shao2024deepseekmathgrpo), researchers have extended it to the multimodal understanding domain, including visual question answering (huang2025visionr1; liu2025visualrft). In image generation, RL has also been shown to be effective (jiang2025t2i-r1; tong2025delving; jiang2025co; pan2025focusdiff; duan2025got-r1; pan2025SelfReflectiveRL; liu2025flowgrpo; xue2025dancegrpo; xiao2025mindomni), but has been used as a training-time optimization method. Unlike them, we employ a simple algorithm, REINFORCE (williams1992simple), for test-time optimization.
@@ 3 Method
Figure 2: Overview of MILR. MILR performs test-time latent reasoning in a unified latent space; it uses policy gradients to iteratively refine text & image latents $\mathbf{z}^{(t)},\mathbf{z}^{(v)}$, guided by a reward model. The reward model scores each generated image conditioned on the initial prompt.
@@ 3.1 Reasoning-augmented multimodal image generation
We are interested in a framework for reasoning-augmented image generation that natively supports language reasoning before image synthesis. A typical implementation is through unified multimodal understanding and generation (MUG; chen2025janusprounifiedmultimodalunderstanding; deng2025bagel). Suppose the language token sequence $\mathbf{t}\coloneqq t_{1:M}\coloneqq t_{1},t_{2},\ldots,t_{M}$, the image token sequence $\mathbf{v}\coloneqq v_{1:N}\coloneqq v_{1},v_{2},\ldots,v_{N}$, and the given instruction $c$, MUG defines multi-modal image generation as an autoregressive generation process:
@@
The generation of $\mathbf{v}$ depends on the reasoning via language tokens $\mathbf{t}$, which are generated from $c$. By analogy to textual reasoning, we refer to the generation of $\mathbf{v}$ as visual reasoning. To produce the final image, $\mathbf{v}$ is further passed through a pixel decoder: $V_{f}\sim p(\cdot|\mathbf{t},\mathbf{v},c)$. Since we focus on discrete image tokens produced by a pre-trained discrete VAE (chen2025janusprounifiedmultimodalunderstanding), the pixel decoder $p(\cdot|\mathbf{t},\mathbf{v},c)$ becomes deterministic given $\mathbf{v}$.
@@ 3.2.1 Multi-Modal Latent Reasoning
Rather than searching over discrete image and text tokens, we propose searching in the unified latent vector space, that is, searching over their continuous vector representations. As in MUG, these vectors correspond to the intermediate model outputs at respective token positions. They lie in a vector space that encodes both image and text tokens, offering a unified view of visual and textual reasoning and facilitating cross-modal reasoning (see Figure 2).
@@ 3.2.2 Gradient-based optimization for latent reasoning
In general, the problem defined by Equation 3 does not admit a closed-form solution, so we resort to REINFORCE (williams1992simple), a policy gradient optimization method. We note that it has been applied to fully textual reasoning in language tasks (li2025seekdarkreasoningtesttime), but we, for the first time, extend it to unified multimodal latent reasoning for image generation.
@@
Naively, we would optimize [^1] We experimented with different optimization strategies, see details in Section A.1. all $M+N$ latents in $z_{1:M+N}$, but searching using only the guidance of a reward model is potentially biased, and it does not leverage the generative capacity of MUG for better exploration; instead, we optimize only the first $\lambda_{t}M$ (where $\lambda_{t}\in(0,1]$) latents for text. After decoding them into discrete tokens, we complete textual reasoning via the standard autoregressive generation conditioned on them. As we will later see, this simple strategy strikes a good balance between efficiency and performance (see Section 4.3.1). For visual reasoning, we adopt a similar strategy and optimize the first $\lambda_{v}N$ (where $\lambda_{v}\in(0,1]$) latents. [^2] In each iteration, the text token number $M$ may vary while the image token number $N$ remains unchanged. This is further supported by the observations of hu2025anchortokenmatchingimplicit: the first few tokens govern the global structure of the image, while the remaining tokens primarily influence high-frequency details. [^3] Rather than optimizing a prefix of image token sequence, we can optimize a random subset of image tokens, but this strategy proves worse than prefix optimization (see Table 5 in Section A.1).
@@ 4.1 Experimental Settings
Benchmarks and Baselines. We conduct experiments on three benchmarks widely used in image generation: GenEval (ghosh2023geneval), T2I-CompBench (huang2023t2icompbench), and WISE (niu2025wise). We compare MILR against the following three sets of baselines:
@@
  • Training-based reasoning models refer to the image generation models that acquire the reasoning ability through training, including GoT-R1 (duan2025got-r1), T2I-R1 (jiang2025t2i-r1), Flow-GRPO (liu2025flowgrpo), and GRPO- and DPO-tuned Janus-Pro (tong2025delving).
@@
  • Test-time reasoning models refer to the models that admit reasoning through tailored inference strategies, including Reflect-DiT (li2025reflectdit) and ReflectionFlow (zhuo2025reflectionflow), which uses language feedback, and Best-of-N and PARM (guo2025can), which rely on search.
@@
Hyperparameters and configurations. We choose Janus-Pro (chen2025janusprounifiedmultimodalunderstanding), an autoregressive model, as the MUG model. For the portions of text and image tokens that are optimized, we perform grid search on a validation split sampled from GenEval and empirically set $\lambda_{t}=0.2$ for text and $\lambda_{v}=0.02$ for images (see Figure 5). We use the Adam optimizer (kingma2014adam), where the learning rate is empirically set to 0.03. For each benchmark, we use its own evaluation toolkit as the reward model, following previous work (liu2025flowgrpo; jiang2025t2i-r1; tong2025delving). We conduct all experiments on a single NVIDIA A100 80GB GPU. A discussion on model efficiency can be found in Section A.3.3.
@@
Table 1: Results on GenEval. The best score is in bold, and the second best is underlined.
@@
Method Single Obj. $\uparrow$ Two Obj. $\uparrow$ Counting $\uparrow$ Colors $\uparrow$ Position $\uparrow$ Attr. Binding $\uparrow$ Overall $\uparrow$ Non-reasoning Models LlamaGen (sun2024LlamaGen) 0.71 0.34 0.21 0.58 0.07 0.04 0.32 Emu3 (wang2024emu3) 0.98 0.71 0.34 0.81 0.17 0.21 0.54 FLUX.1-dev (flux2024) 0.98 0.79 0.73 0.77 0.22 0.45 0.66 DALL-E 3 (betker2023dalle) 0.96 0.87 0.47 0.83 0.43 0.45 0.67 SD3-Medium (esser2024sd3-med) 0.99 0.94 0.72 0.89 0.33 0.60 0.74 BAGEL (deng2025bagel) 0.99 0.94 0.81 0.88 0.64 0.63 0.82 GPT-4o (openai_4o_image_generation) 0.99 0.92 0.85 0.91 0.75 0.66 0.85 Training-based Reasoning Models GoT-R1 (duan2025got-r1) 0.99 0.94 0.50 0.90 0.46 0.68 0.75 T2I-R1 (jiang2025t2i-r1) 0.99 0.91 0.53 0.91 0.76 0.65 0.79 Flow-GRPO (liu2025flowgrpo) 1.00 0.99 0.95 0.92 0.99 0.86 0.95 ReasonGen-R1 (zhang2025reasongenr1cotautoregressiveimage) 0.99 0.94 0.62 0.90 0.84 0.84 0.86 Janus-Pro-7B(+GRPO) (tong2025delving) 0.99 0.87 0.61 0.87 0.82 0.68 0.81 Janus-Pro-7B(+DPO) (guo2025can) 0.99 0.89 0.65 0.92 0.82 0.72 0.83 Test-time Reasoning Models Reflect-DiT (li2025reflectdit) 0.98 0.96 0.80 0.88 0.66 0.60 0.81 ReflectionFlow (zhuo2025reflectionflow) 1.00 0.98 0.90 0.96 0.93 0.72 0.91 Janus-Pro-7B(+Text Enhanced Reasoning) 0.98 0.91 0.55 0.89 0.74 0.67 0.79 Janus-Pro-7B(+Best-of-N) 0.99 0.96 0.89 0.93 0.92 0.80 0.91 Janus-Pro-7B(+PARM) (guo2025can) 1.00 0.95 0.80 0.93 0.91 0.85 0.91 Janus-Pro-1B (chen2025janusprounifiedmultimodalunderstanding) 0.98 0.82 0.51 0.89 0.65 0.56 0.73 Janus-Pro-1B+MILR 1.00 0.91 0.78 0.92 0.86 0.86 0.89 Janus-Pro 7B (chen2025janusprounifiedmultimodalunderstanding) 0.98 0.85 0.56 0.89 0.77 0.64 0.78 Janus-Pro-7B+MILR 1.00 0.96 0.90 0.98 0.98 0.91 0.95
@@
Method T2I-CompBench WISE Color $\uparrow$ Shape $\uparrow$ Texture $\uparrow$ Spatial $\uparrow$ Non-Spatial $\uparrow$ Complex. $\uparrow$ Overall $\uparrow$ Avg $\uparrow$ Non-reasoning Models PixArt-$\alpha$ chen2023pixart 0.6690 0.4927 0.6477 0.2064 0.3197 0.3433 0.4465 0.47 FLUX.1-dev flux2024 0.7407 0.5718 0.6922 0.2863 0.3127 0.3703 0.4957 0.50 DALL-E 3 betker2023dalle 0.7785 0.6209 0.7036 0.2865 0.3003 0.3773 0.5112 - SD3-Medium esser2024sd3-med 0.8132 0.5885 0.7334 0.3200 0.3140 0.3771 0.5244 0.42 Show-o xie2024showo 0.5600 0.4100 0.4600 0.2000 0.3000 0.2900 0.3700 0.30 BAGEL deng2025bagel 0.8027 0.5685 0.7021 0.3488 0.3101 0.3824 0.5191 0.52 Training-based Reasoning Models T2I-R1 jiang2025t2i-r1 0.8130 0.5852 0.7243 0.3378 0.3090 0.3993 0.5281 0.54 GoT-R1 duan2025got-r1 0.8139 0.5549 0.7339 0.3306 0.3169 0.3944 0.5241 - Janus-Pro-7B(+GRPO) tong2025delving 0.7721 0.5366 0.7317 0.2869 0.3087 0.3697 0.5010 - Test-time Reasoning Models Show-o + PARM guo2025can 0.7500 0.5600 0.6600 0.2900 0.3100 0.3700 0.4900 - Janus-Pro-7B(+Text Enhanced Reasoning) 0.7087 0.4419 0.5821 0.2597 0.3072 0.3761 0.4459 0.46 Janus-Pro-7B(+Best-of-N) 0.7089 0.4925 0.7089 0.3542 0.3262 0.3721 0.4938 0.52 Janus-Pro-1B chen2025janusprounifiedmultimodalunderstanding 0.3411 0.2261 0.2696 0.0968 0.2808 0.2721 0.2478 0.26 Janus-Pro-1B+MILR 0.6066 0.2796 0.4177 0.2796 0.2622 0.2613 0.3512 0.40 Janus-Pro-7B chen2025janusprounifiedmultimodalunderstanding 0.6359 0.3528 0.4936 0.2061 0.3085 0.3559 0.3921 0.35 Janus-Pro-7B+MILR 0.8508 0.5117 0.6949 0.4613 0.3078 0.3684 0.5325 0.63

### v3

@@ 1 Introduction
Text-guided image generation is the task of synthesizing an image conditioned on a given text instruction. In recent years, the field has witnessed transformative progress: moving from generative adversarial models (Goodfellow et al., 2014) to autoregressive and diffusion approaches (Sun et al., 2024; Esser et al., 2024; Black Forest Labs, 2024; Chen et al., 2025; Wu et al., 2025a). However, traditional models are limited in generating images in a single-shot fashion and thus are unable to resolve potential defects (Huang et al., 2025a; Niu et al., 2025). Inspired by the success of reasoning-augmented LLMs—such as OpenAI o1 (OpenAI et al., 2024) and DeepSeek-R1 (DeepSeek-AI et al., 2025)—that can reflect on and refine their thoughts, and answer accordingly, recent works have attempted to endow image generation models with reasoning ability (Guo et al., 2025; Fang et al., 2025; Zhang et al., 2025b; Wu et al., 2025b; Chern et al., 2025).
Reasoning-augmented image generation models typically perform reasoning in two spaces: language and image spaces. Language reasoning primarily involves refining instructions to make them more comprehensible to the model (Wu et al., 2025b; Li et al., 2025b), and image reasoning iterates on the generation guided by a quality metric (Guo et al., 2025; Zhuo et al., 2025). Early works implement reasoning either in the language or in the image space, lacking a mechanism for synergistic reasoning across the two spaces. To fill the gap, recent works resort to unified multimodal understanding and generation (MUG; Jiang et al. (2025a); Duan et al. (2025); Zhang et al. (2025b)), which natively supports language reasoning before generating images (referred to as multimodal image generation), facilitating cross-modal image-text reasoning. Despite the success, these approaches require carefully curated reasoning data and depend on model fine-tuning, rendering them complex and costly to develop in practice.
@@
Reasoning with MILR involves finding the best vector representations of image and text tokens that lead to improved image quality. We implement it using the policy gradient method (Williams, 1992), where, at test time, the reward is computed by scoring the compatibility between the generated image and the given instruction. Crucially, gradients are only back-propagated to the cross-modal latent representations (i.e., the intermediate model outputs) obtained from the forward pass (see Figure 1), without altering any model parameters, and thus making MILR a test-time latent reasoning method.
We evaluate MILR on three widely-used image generation benchmarks: GenEval (Ghosh et al., 2023), T2I-CompBench (Huang et al., 2023), and WISE (Niu et al., 2025). MILR established new state-of-the-art results across all benchmarks. Notably, it achieves an overall score of 0.95 on GenEval, matching the best training-based model and outperforming the best test-time-scaling method by 4.4%. On the more challenging WISE benchmark, MILR obtains an overall score of 0.63, surpassing the strongest baseline model by 16.7%. Our further analysis reveals that the best performance of MILR is driven by its ability to perform joint image-text reasoning in a unified latent space. This unique advantage also makes MILR successful in instructions that involve challenging cultural and temporal reasoning.
@@ 2 Related Works
Remarkable progress has been achieved in reasoning-augmented LLMs (Wei et al., 2022; DeepSeek-AI et al., 2025), but effectively integrating reasoning into conventional text-guided image generation models remains a challenge (Betker et al., 2023; Sun et al., 2024; Esser et al., 2024; Zhao et al., 2024; Chen et al., 2025). With the development of unified multimodal understanding and generation that can generate images and text in an interleaved manner, many works have explored reasoning for text-guided image generation. The predominant paradigm uses causal language modeling to unify image and text generation as next token prediction, enabling the model to make plans via chain-of-thought before generating images (Fang et al., 2025; Deng et al., 2025; Xiao et al., 2025; Cai et al., 2025; Guo et al., 2025; Jiang et al., 2025a; Duan et al., 2025). Another line of research focuses on reasoning over images via the test-time scaling strategy (Li et al., 2025b; Zhuo et al., 2025). These works typically rely on an external critic model to provide feedback on further improvement. Unlike all these methods that explicitly perform reasoning over raw images and text, we use test-time optimization to refine the latent representations of image and text tokens, leading to a unified cross-modal latent reasoning method.
@@
Differently from explicit reasoning via a chain of thoughts (Wei et al., 2022), latent reasoning refers to an implicit reasoning mechanism applied to the latent states (e.g., intermediate outputs) of the model. As in Transformer-based models, latent reasoning is typically implemented as spatial and temporal recurrences. Spatial recurrences implicitly deepen the model by iterating over the latent states between the Transformer layers (Hao et al., 2024; Cheng and Van Durme, 2024; Zhang et al., 2025a; Shen et al., 2025), while temporal recurrences refine the latent states through iterations across input tokens (Dao and Gu, 2024; Geiping et al., 2025). Under the recurrences is the idea of scaling up test-time computation for inference; however, this requires the recurrent modules to be pre-trained. In a similar vein to those, we iterate over the unified image-text latent states only at test time, without introducing and updating any model parameters.
@@
Reinforcement learning (RL) has been the key to eliciting the reasoning ability of large language models (DeepSeek-AI et al., 2025; OpenAI et al., 2024). Inspired by the success of GRPO, which is an improved RL algorithm over PPO (Schulman et al., 2017) and is used in DeepSeek-R1 (Shao et al., 2024), researchers have extended it to the multimodal understanding domain, including visual question answering (Huang et al., 2025b; Liu et al., 2025b). In image generation, RL has also been shown to be effective (Jiang et al., 2025a; Tong et al., 2025; Jiang et al., 2025b; Pan et al., 2025b; Duan et al., 2025; Pan et al., 2025a; Liu et al., 2025a; Xue et al., 2025; Xiao et al., 2025), but has been used as a training-time optimization method. Unlike them, we employ a simple algorithm, REINFORCE (Williams, 1992), for test-time optimization.
@@ 3 Method
Figure 2: Overview of MILR. MILR performs test-time latent reasoning in a unified latent space; it uses policy gradients to iteratively refine text & image latents $\mathbf{z}^{(t)},\mathbf{z}^{(v)}$, guided by a reward model. The reward model scores each generated image conditioning on the instruction.
@@ 3.1 Reasoning-augmented multimodal image generation
We are interested in a framework for reasoning-augmented image generation that natively supports language reasoning before image synthesis. A typical implementation is through unified multimodal understanding and generation (MUG; Chen et al. (2025); Deng et al. (2025)). Suppose the language token sequence $\mathbf{t}\coloneqq t_{1:M}\coloneqq t_{1},t_{2},\ldots,t_{M}$, the image token sequence $\mathbf{v}\coloneqq v_{1:N}\coloneqq v_{1},v_{2},\ldots,v_{N}$, and the given instruction $c$, MUG defines multi-modal image generation as an autoregressive generation process:
@@
The generation of $\mathbf{v}$ depends on the reasoning via language tokens $\mathbf{t}$, which are generated from $c$. By analogy to textual reasoning, we refer to the generation of $\mathbf{v}$ as visual reasoning. To produce the final image, $\mathbf{v}$ is further passed through a pixel decoder: $V_{f}\sim p(\cdot|\mathbf{t},\mathbf{v},c)$. Since we focus on discrete image tokens produced by a pre-trained discrete VAE (Chen et al., 2025), the pixel decoder $p(\cdot|\mathbf{t},\mathbf{v},c)$ becomes deterministic given $\mathbf{v}$.
@@ 3.2.1 Multi-Modal Latent Reasoning
Rather than searching over discrete image and text tokens, we propose searching in the unified latent vector space, that is, searching over their continuous vector representations. As in MUG, these vectors correspond to the intermediate model outputs at respective token positions. They lie in a vector space that encodes both image and text tokens, offering a unified view of visual and textual reasoning [^1] We refer to “reasoning” as iterative updates in the latent space rather than text-based chain-of-thought. and facilitating cross-modal reasoning (see Figure 2).
@@ 3.2.2 Gradient-based optimization for latent reasoning
In general, the problem defined by Equation 3 does not admit a closed-form solution, so we resort to REINFORCE (Williams, 1992), a policy gradient optimization method. We note that it has been applied to fully textual reasoning in language tasks (Li et al., 2025a), but we, for the first time, extend it to unified multimodal latent reasoning for image generation.
@@
Naively, we would optimize [^2] We experimented with different optimization strategies, see details in Section A.1. all $M+N$ latents in $z_{1:M+N}$, but searching using only the guidance of a reward model is potentially biased, and it does not leverage the generative capacity of MUG for better exploration; instead, we optimize only the first $\lambda_{t}M$ (where $\lambda_{t}\in(0,1]$) latents for text. After decoding them into discrete tokens, we complete textual reasoning via the standard autoregressive generation conditioned on them. As we will later see, this simple strategy strikes a good balance between efficiency and performance (see Section 4.3.1). For visual reasoning, we adopt a similar strategy and optimize the first $\lambda_{v}N$ (where $\lambda_{v}\in(0,1]$) latents. [^3] In each iteration, the text token number $M$ may vary while the image token number $N$ remains unchanged. This is further supported by the observations of Hu et al. (2025): the first few tokens govern the global structure of the image, while the remaining tokens primarily influence high-frequency details. [^4] Rather than optimizing a prefix of image token sequence, we can optimize a random subset of image tokens, but this strategy proves worse than prefix optimization (see Table 5 in Section A.1).
@@ 4.1 Experimental Settings
Benchmarks and Baselines. We conduct experiments on three benchmarks widely used in image generation: GenEval (Ghosh et al., 2023), T2I-CompBench (Huang et al., 2023), and WISE (Niu et al., 2025). We compare MILR against the following three sets of baselines:
@@
  • Training-based reasoning models refer to the image generation models that acquire the reasoning ability through training, including GoT-R1 (Duan et al., 2025), T2I-R1 (Jiang et al., 2025a), Flow-GRPO (Liu et al., 2025a), and GRPO- and DPO-tuned Janus-Pro (Tong et al., 2025).
@@
  • Test-time reasoning models refer to the models that admit reasoning through tailored inference strategies, including Reflect-DiT (Li et al., 2025b) and ReflectionFlow (Zhuo et al., 2025), which uses language feedback, and Best-of-N and PARM (Guo et al., 2025), which rely on search.
@@
Hyperparameters and configurations. We choose Janus-Pro (Chen et al., 2025), an autoregressive model, as the MUG model. For the portions of text and image tokens that are optimized, we perform grid search on a validation split sampled from GenEval and empirically set $\lambda_{t}=0.2$ for text and $\lambda_{v}=0.02$ for images (see Figure 5). We use the Adam optimizer (Kingma, 2014), where the learning rate is empirically set to 0.03. For each benchmark, we use its own evaluation toolkit as the reward model, following previous work (Liu et al., 2025a; Jiang et al., 2025a; Tong et al., 2025). We conduct all experiments on a single NVIDIA A100 80GB GPU. A discussion on model efficiency can be found in Section A.3.3.
@@
Table 1: Results on GenEval. The best score is in bold and the second best is underlined.
@@
Method Single Obj. $\uparrow$ Two Obj. $\uparrow$ Counting $\uparrow$ Colors $\uparrow$ Position $\uparrow$ Attr. Binding $\uparrow$ Overall $\uparrow$ Non-reasoning Models LlamaGen (Sun et al., 2024) 0.71 0.34 0.21 0.58 0.07 0.04 0.32 Emu3 (Wang et al., 2024) 0.98 0.71 0.34 0.81 0.17 0.21 0.54 FLUX.1-dev (Black Forest Labs, 2024) 0.98 0.79 0.73 0.77 0.22 0.45 0.66 DALL-E 3 (Betker et al., 2023) 0.96 0.87 0.47 0.83 0.43 0.45 0.67 SD3-Medium (Esser et al., 2024) 0.99 0.94 0.72 0.89 0.33 0.60 0.74 BAGEL (Deng et al., 2025) 0.99 0.94 0.81 0.88 0.64 0.63 0.82 GPT-4o (OpenAI, 2025) 0.99 0.92 0.85 0.91 0.75 0.66 0.85 Training-based Reasoning Models GoT-R1 (Duan et al., 2025) 0.99 0.94 0.50 0.90 0.46 0.68 0.75 T2I-R1 (Jiang et al., 2025a) 0.99 0.91 0.53 0.91 0.76 0.65 0.79 Flow-GRPO (Liu et al., 2025a) 1.00 0.99 0.95 0.92 0.99 0.86 0.95 ReasonGen-R1 (Zhang et al., 2025b) 0.99 0.94 0.62 0.90 0.84 0.84 0.86 Janus-Pro-7B(+GRPO) (Tong et al., 2025) 0.99 0.87 0.61 0.87 0.82 0.68 0.81 Janus-Pro-7B(+DPO) (Guo et al., 2025) 0.99 0.89 0.65 0.92 0.82 0.72 0.83 Test-time Reasoning Models Reflect-DiT (Li et al., 2025b) 0.98 0.96 0.80 0.88 0.66 0.60 0.81 ReflectionFlow (Zhuo et al., 2025) 1.00 0.98 0.90 0.96 0.93 0.72 0.91 Janus-Pro-7B(+Text Enhanced Reasoning) 0.98 0.91 0.55 0.89 0.74 0.67 0.79 Janus-Pro-7B(+Best-of-N) 0.99 0.96 0.89 0.93 0.92 0.80 0.91 Janus-Pro-7B(+PARM) (Guo et al., 2025) 1.00 0.95 0.80 0.93 0.91 0.85 0.91 Janus-Pro-1B (Chen et al., 2025) 0.98 0.82 0.51 0.89 0.65 0.56 0.73 Janus-Pro-1B+MILR 1.00 0.91 0.78 0.92 0.86 0.86 0.89 Janus-Pro 7B (Chen et al., 2025) 0.98 0.85 0.56 0.89 0.77 0.64 0.78 Janus-Pro-7B+MILR 1.00 0.96 0.90 0.98 0.98 0.91 0.95
@@
Method T2I-CompBench WISE Color $\uparrow$ Shape $\uparrow$ Texture $\uparrow$ Spatial $\uparrow$ Non-Spatial $\uparrow$ Complex. $\uparrow$ Overall $\uparrow$ Avg $\uparrow$ Non-reasoning Models PixArt-$\alpha$ Chen et al. (2023) 0.6690 0.4927 0.6477 0.2064 0.3197 0.3433 0.4465 0.47 FLUX.1-dev Black Forest Labs (2024) 0.7407 0.5718 0.6922 0.2863 0.3127 0.3703 0.4957 0.50 DALL-E 3 Betker et al. (2023) 0.7785 0.6209 0.7036 0.2865 0.3003 0.3773 0.5112 - SD3-Medium Esser et al. (2024) 0.8132 0.5885 0.7334 0.3200 0.3140 0.3771 0.5244 0.42 Show-o Xie et al. (2024) 0.5600 0.4100 0.4600 0.2000 0.3000 0.2900 0.3700 0.30 BAGEL Deng et al. (2025) 0.8027 0.5685 0.7021 0.3488 0.3101 0.3824 0.5191 0.52 Training-based Reasoning Models T2I-R1 Jiang et al. (2025a) 0.8130 0.5852 0.7243 0.3378 0.3090 0.3993 0.5281 0.54 GoT-R1 Duan et al. (2025) 0.8139 0.5549 0.7339 0.3306 0.3169 0.3944 0.5241 - Janus-Pro-7B(+GRPO) Tong et al. (2025) 0.7721 0.5366 0.7317 0.2869 0.3087 0.3697 0.5010 - Test-time Reasoning Models Show-o + PARM Guo et al. (2025) 0.7500 0.5600 0.6600 0.2900 0.3100 0.3700 0.4900 - Janus-Pro-7B(+Text Enhanced Reasoning) 0.7087 0.4419 0.5821 0.2597 0.3072 0.3761 0.4459 0.46 Janus-Pro-7B(+Best-of-N) 0.7089 0.4925 0.7089 0.3542 0.3262 0.3721 0.4938 0.52 Janus-Pro-1B Chen et al. (2025) 0.3411 0.2261 0.2696 0.0968 0.2808 0.2721 0.2478 0.26 Janus-Pro-1B+MILR 0.6066 0.2796 0.4177 0.2796 0.2622 0.2613 0.3512 0.40 Janus-Pro-7B Chen et al. (2025) 0.6359 0.3528 0.4936 0.2061 0.3085 0.3559 0.3921 0.35 Janus-Pro-7B+MILR 0.8508 0.5117 0.6949 0.4613 0.3078 0.3684 0.5325 0.63